OpenStudy (anonymous):

Rationalize the denominator of (√18)/(√3) 6 (2√3)/3 (3√3)/2 √6

OpenStudy (anonymous):

D option

OpenStudy (ash2326):

@Bferg18 I'll explain you \[\frac{\sqrt{18}}{\sqrt 3}\] Multiply numerator and denominator by \(\sqrt 3\) \[\frac{\sqrt{18}}{\sqrt 3}\times \frac{\sqrt 3}{\sqrt 3}\] What would you get?